# Completeness of coalition logics with seriality, independence of agents, or determinism

### v1

@@ 3.1 Our approach
The basic step of the induction can be directly done by the classical propositional logic, and the inductive step is crucial. We will go through the inductive step in the following way.
@@
Second, we show the downward validity lemma: for every standard disjunction $\phi$, if $\phi$ is valid, then the validity-reduction-condition of $\phi$ is met. Here the validity-reduction-condition of $\phi$ concerns the validity of some formulas with lower modal depth than $\phi$.
Third, we show the upward derivability lemma: for every standard disjunction $\phi$, if the derivability-reduction-condition of $\phi$ is met, then $\phi$ is derivable. Here, the derivability-reduction-condition of $\phi$ is the result of replacing validity in the validity-reduction-condition of $\phi$ by derivability.
Note a conjunction is valid (derivable) if and only if all of its conjuncts are valid (derivable). Thus, the two lemmas would let us go through the inductive step. Note by soundness and completeness, the other direction of the two lemmas also holds, which is why their names contain “reduction”.
The downward validity lemma is the most crucial step for the whole proof. We will show it by showing its contrapositive, called the upward satisfiability lemma: for every standard conjunction $\phi^{\prime}$, if the satisfiability-reduction-condition of $\phi^{\prime}$ is met, then $\phi^{\prime}$ is satisfiable. Here, a standard conjunction $\phi^{\prime}$ is equivalent to the negation of a standard disjunction $\phi$, and the satisfiability-reduction-condition of $\phi^{\prime}$ is equivalent to the negation of the validity-reduction-condition of $\phi$.
To show the upward satisfiability lemma, the key is to find a pointed model satisfying a standard conjunction $\phi^{\prime}$, given its satisfiability-reduction-condition is met. To achieve this, we will use a method called the blueprint method.
3.2 Standard formulas and a normal form lemma
For every $n\in\mathbb{N}$, the set $\{-x\mid x\in\mathbb{N}\text{ and }1\leq x\leq n\}$ is called a negative indice set and the set $\{x\mid x\in\mathbb{N}\text{ and }0\leq x\leq n\}$ is called a positive indice set. Note negative indice sets are allowed to be empty, but positive indice sets are not.
@@ 3.2 Standard formulas and a normal form lemma
A standard disjunction is a formula in the form $\gamma\vee(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\to\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\phi_{j})$, where $\gamma$ is an elementary disjunction, $\mathbb{I}$ is a negative indice set, and $\mathbb{J}$ is a positive indice set such that (1) if $\mathbb{I}\neq\emptyset$, then $\langle[\mathrm{A}_{i}]\rangle\phi_{i}=\langle[\emptyset]\rangle\top$ for some $i\in\mathbb{I}$, and (2) $\langle[\mathrm{B}_{0}]\rangle\phi_{0}=\langle[\mathtt{AG}]\rangle\bot$.
A standard conjunction is a formula in the form $\gamma\wedge\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\wedge\bigwedge_{j\in\mathbb{J}}\neg\langle[\mathrm{B}_{j}]\rangle\phi_{j}$, where $\gamma$ is an elementary conjunction, $\mathbb{I}$ is a negative indice set, and $\mathbb{J}$ is a positive indice set such that (1) if $\mathbb{I}\neq\emptyset$, then $\langle[\mathrm{A}_{i}]\rangle\phi_{i}=\langle[\emptyset]\rangle\top$ for some $i\in\mathbb{I}$, and (2) $\neg\langle[\mathrm{B}_{0}]\rangle\phi_{0}=\neg\langle[\mathtt{AG}]\rangle\bot$.
@@
We define the modal depth of formulas as usual.
@@
Let $\mathsf{X}\in\mathrm{ES}$. For every $\phi\in\Phi$, there is $\phi^{\prime}$ such that (1) $\vdash_{\mathsf{MCL}+\mathsf{X}}\phi\leftrightarrow\phi^{\prime}$, (2) $\phi$ and $\phi^{\prime}$ have the same modal depth, and (3) $\phi^{\prime}$ is in the form of $\chi_{0}\land\dots\land\chi_{k}$, where every $\chi_{i}$ is a standard disjunction.
[…]
The following notions will be used in stating the downward validity lemma and the upward derivability lemma.
Definition 9 (Sets of basic positive indices).
Let $\gamma\vee(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\to\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$ be a standard disjunction. $\mathbb{J}_{0}=\{j\in\mathbb{J}\mid\mathrm{B}_{i}=\mathtt{AG}\}$ is called the set of basic positive indices.
Definition 10 (Neat sets of negative indices).
Let $\mathsf{X}\in\mathrm{ES}$ and $\gamma\vee(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\to\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$ be a standard disjunction.
  1. (1)
  2. (2)
    if $\mathsf{S}\notin\mathsf{X}$, then $\mathbb{I}^{\prime}\neq\emptyset$;
  3. (3)
The following fact can be easily shown and we skip its proof:
@@
Let $\gamma\vee(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\to\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$ be a standard disjunction. Then, for all $\mathsf{X}\in\mathrm{ES}$ and $\mathbb{I}^{\prime},\mathbb{I}^{\prime\prime}\subseteq\mathbb{I}$, if $\mathbb{I}^{\prime}\subseteq\mathbb{I}^{\prime\prime}$ and $\mathbb{I}^{\prime\prime}$ is $\mathsf{X}$-neat, then $\mathbb{I}^{\prime}$ is $\mathsf{X}$-neat.
Roughly, the point of the fact is that neatness is conversely monotonic for negative indice sets.
3.3 Downward validity lemma
3.3.1 Blueprints and their realization
@@
A blueprint is a tuple $\mathtt{BP}=(\mathtt{AG},\mathtt{AC},\mathtt{list}_{\mathtt{AG}})$, where:
  • $\mathtt{AG}$ is a finite set of agents;
@@
For any $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, intuitively, $\mathtt{list}_{\mathtt{AG}}(\sigma_{\mathtt{AG}})$ lists some formulas which $\sigma_{\mathtt{AG}}$ enables.
Definition 12 (Performable joint actions).
Let $\mathtt{BP}=(\mathtt{AG},\mathtt{AC},\mathtt{list}_{\mathtt{AG}})$ be a blueprint.
@@
Let $\mathtt{BP}=(\mathtt{AG},\mathtt{AC},\mathtt{list}_{\mathtt{AG}})$ be a blueprint and $\mathsf{X}\in\mathrm{ES}$.
@@
Let $\mathsf{X}\in\mathrm{ES}$ and $\mathtt{BP}=(\mathtt{AG},\mathtt{AC},\mathtt{list}_{\mathtt{AG}})$ be an $\mathsf{X}$-regular blueprint.
@@
Let $\mathsf{X}\in\mathrm{ES}$, $\mathtt{BP}=(\mathtt{AG},\mathtt{AC}_{0},\mathtt{list}_{\mathtt{AG}})$ be an $\mathsf{X}$-regular blueprint, and $\gamma$ be an $\mathsf{X}$-satisfiable elementary conjunction.
We say a pointed $\mathsf{X}$-model $(\mathtt{M},s_{0})$, where $\mathtt{M}=(\mathtt{ST},\mathtt{AG},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathtt{AP},\mathsf{label})$, realizes $\mathtt{BP}$ and $\gamma$ if the following conditions are met:
@@
Let $\mathtt{BP}=(\mathtt{AG},\mathtt{AC}_{0},\mathtt{list}_{\mathtt{AG}})$ be an $\mathsf{X}$-regular blueprint, and $\gamma$ be an $\mathsf{X}$-satisfiable elementary conjunction.
Let $\{(\mathtt{M}^{\lambda}_{\phi},s^{\lambda}_{\phi})\mid\lambda\in\mathtt{PJA}_{\mathtt{AG}}\text{ and }\phi\in\mathtt{list}_{\mathtt{AG}}(\lambda)\}$, where $\mathtt{M}^{\lambda}_{\phi}=(\mathtt{ST}^{\lambda}_{\phi},\mathtt{AG},\mathtt{AC}^{\lambda}_{\phi},{\mathsf{out}_{\mathtt{AG}}}^{\lambda}_{\phi},\mathtt{AP},\mathsf{label}^{\lambda}_{\phi})$, be a set of pointed $\mathsf{X}$-models meeting the following conditions:
  1. (1)
@@
Let $s_{0}$ be a state not in any of these models.
Define a pointed model $(\mathtt{M},s_{0})$, where $\mathtt{M}=(\mathtt{ST},\mathtt{AG},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathtt{AP},\mathsf{label})$, as follows:
@@
    ${\mathsf{out}_{\mathtt{AG}}(s,{\sigma_{\mathtt{AG}}})}=\begin{cases}{{\mathsf{out}_{\mathtt{AG}}}^{\lambda}_{\phi}(s,{\sigma_{\mathtt{AG}}}})&\text{if }s\in{\mathtt{ST}^{\lambda}_{\phi}}\text{ and }{\sigma_{\mathtt{AG}}}\in{\mathtt{JA}_{\mathtt{AG}}}^{\lambda}_{\phi}\\ &\text{for some }\lambda\in\mathtt{PJA}_{\mathtt{AG}}\text{ and }\phi\in\mathtt{list}_{\mathtt{AG}}(\lambda)\\ \{s^{{\sigma_{\mathtt{AG}}}}_{\phi}\mid\phi\in\mathtt{list}_{\mathtt{AG}}({\sigma_{\mathtt{AG}}})\}&\text{if }s=s_{0}\text{ and }{\sigma_{\mathtt{AG}}}\in\mathtt{PJA}_{\mathtt{AG}}\\ \emptyset&\text{otherwise}\end{cases}$
@@
    ${\mathsf{label}(s)}=\begin{cases}{\mathsf{label}^{\lambda}_{\phi}(s)}&\text{if}s\in{\mathtt{ST}^{\lambda}_{\phi}}\text{ for some }\lambda\in\mathtt{PJA}_{\mathtt{AG}}\text{ and }\phi\in\mathtt{list}_{\mathtt{AG}}(\lambda)\\ \{p\mid\text{$p$ is a conjunct of $\gamma$}\}&\text{if }s=s_{0}\\ \end{cases}$
@@
Now we show the last three conditions given in the definition of realization of blueprints hold.
@@
    Let $s\in\mathsf{out}_{\mathrm{C}}(s_{0},\sigma_{\mathrm{C}})$. Then $s\in\mathsf{out}_{\mathtt{AG}}(s_{0},\sigma_{\mathtt{AG}})$ for some $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathtt{AG}}\subseteq\sigma_{\mathrm{C}}$. Then, $s=s^{\sigma_{\mathtt{AG}}}_{\phi}$ for some $\phi\in\mathtt{list}_{\mathtt{AG}}(\sigma_{\mathtt{AG}})$. Note $\mathtt{M}^{\sigma_{\mathtt{AG}}}_{\phi},s^{\sigma_{\mathtt{AG}}}_{\phi}\Vdash\phi$. Then $\mathtt{M},s^{\sigma_{\mathtt{AG}}}_{\phi}\Vdash\phi$. Then $\mathtt{M},s^{\sigma_{\mathtt{AG}}}_{\phi}\Vdash\psi$, that is, $\mathtt{M},s\Vdash\psi$. Therefore, ${\sigma_{\mathrm{C}}}\vartriangleright_{(\mathtt{M},s_{0})}\psi$.

### v2

@@ 3.1 Our approach: the reduction method
The basic step of the induction can be directly done by the classical propositional logic. We will go through the inductive step as follows.
@@
Second, we show the downward validity lemma: for every standard disjunction $\phi$, if $\phi$ is valid, then the validity-reduction condition of $\phi$ is met. Here, the validity-reduction condition of $\phi$ concerns the validity of some formulas with lower modal depth than $\phi$.
Third, we show the upward derivability lemma: for every standard disjunction $\phi$, if the derivability-reduction condition of $\phi$ is met, then $\phi$ is derivable. Here, the derivability-reduction condition of $\phi$ is the result of replacing validity in the validity-reduction condition of $\phi$ by derivability.
We will show the downward validity lemma by showing its contrapositive, called the upward satisfiability lemma: for every standard conjunction $\phi^{\prime}$, if the satisfiability-reduction condition of $\phi^{\prime}$ is met, then $\phi^{\prime}$ is satisfiable. Here, a standard conjunction $\phi^{\prime}$ is equivalent to the negation of a standard disjunction $\phi$, and the satisfiability-reduction condition of $\phi^{\prime}$ is equivalent to the negation of the validity-reduction condition of $\phi$.
To show the upward satisfiability lemma, the key is to find a pointed model satisfying a standard conjunction $\phi^{\prime}$, given that its satisfiability-reduction condition is met. To achieve this, we will use an important notion, called blueprints.
Remarks
Suppose we have proved that a sound logic is complete by this approach. It is easy to see that the other direction of the downward validity lemma and the upward derivability lemma also holds:
  • For every standard disjunction $\phi$, if the validity-reduction condition of $\phi$ is met, then $\phi$ is valid.
  • For every standard disjunction $\phi$, if $\phi$ is derivable, then the derivability-reduction condition of $\phi$ is met.
This is why we use “reduction” in the names of the two lemmas.
For the reduction method, the downward validity lemma is crucial and other steps are relatively easy.
3.2 A normal form lemma
For every $n\in\mathbb{N}$, the set $\{-x\mid x\in\mathbb{N}\text{ and }1\leq x\leq n\}$ is called a negative indice set and the set $\{x\mid x\in\mathbb{N}\text{ and }0\leq x\leq n\}$ is called a positive indice set. Note that negative index sets are allowed to be empty, but positive index sets are not.
@@ 3.2 A normal form lemma
A standard disjunction is a formula in the form $\gamma\vee(\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\to\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\phi_{j})$, where $\gamma$ is an elementary disjunction, $\mathbb{I}$ is a negative indice set, and $\mathbb{J}$ is a positive indice set such that (1) if $\mathbb{I}\neq\emptyset$, then $\langle\hskip-2.0pt[\mathrm{A}_{-1}]\hskip-2.0pt\rangle\phi_{-1}=\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\top$, and (2) $\langle\hskip-2.0pt[\mathrm{B}_{0}]\hskip-2.0pt\rangle\phi_{0}=\langle\hskip-2.0pt[\mathtt{AG}]\hskip-2.0pt\rangle\bot$.
A standard conjunction is a formula in the form $\gamma\wedge\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\wedge\bigwedge_{j\in\mathbb{J}}\neg\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\phi_{j}$, where $\gamma$ is an elementary conjunction, $\mathbb{I}$ is a negative indice set, and $\mathbb{J}$ is a positive indice set such that (1) if $\mathbb{I}\neq\emptyset$, then $\langle\hskip-2.0pt[\mathrm{A}_{-1}]\hskip-2.0pt\rangle\phi_{-1}=\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\top$, and (2) $\neg\langle\hskip-2.0pt[\mathrm{B}_{0}]\hskip-2.0pt\rangle\phi_{0}=\neg\langle\hskip-2.0pt[\mathtt{AG}]\hskip-2.0pt\rangle\bot$.
@@
The special structure of standard disjunctions and conjunctions will make the statement of the downward validity lemma and the upward derivability lemma simpler.
Example 1 (Standard disjunctions and conjunctions).
Assume $\mathtt{AG}=\{a,b\}$. Let $\gamma=\bot$, $\mathbb{I}=\{-1,-2,-3\}$ and $\mathbb{J}=\{0,1\}$. Then,
$\bot\lor\big{(}\big{(}\langle\hskip-2.0pt[\mathrm{A}_{-1}]\hskip-2.0pt\rangle\phi_{-1}\land\langle\hskip-2.0pt[\mathrm{A}_{-2}]\hskip-2.0pt\rangle\phi_{-2}\land\langle\hskip-2.0pt[\mathrm{A}_{-3}]\hskip-2.0pt\rangle\phi_{-3}\big{)}\rightarrow\big{(}\langle\hskip-2.0pt[\mathrm{B}_{0}]\hskip-2.0pt\rangle\psi_{0}\lor\langle\hskip-2.0pt[\mathrm{B}_{1}]\hskip-2.0pt\rangle\psi_{1}\big{)}\big{)}$
is a standard disjunction, and
$\neg\bot\land\langle\hskip-2.0pt[\mathrm{A}_{-1}]\hskip-2.0pt\rangle\phi_{-1}\land\langle\hskip-2.0pt[\mathrm{A}_{-2}]\hskip-2.0pt\rangle\phi_{-2}\land\langle\hskip-2.0pt[\mathrm{A}_{-3}]\hskip-2.0pt\rangle\phi_{-3}\land\neg\langle\hskip-2.0pt[\mathrm{B}_{0}]\hskip-2.0pt\rangle\psi_{0}\land\neg\langle\hskip-2.0pt[\mathrm{B}_{1}]\hskip-2.0pt\rangle\psi_{1}$
is a standard conjunction, with respect to $\gamma$, $\mathbb{I}$ and $\mathbb{J}$, where $\mathrm{A}_{-1}=\emptyset$, $\phi_{-1}=\top$, $\mathrm{A}_{-2}=\{b\}$, $\phi_{-2}=q$, $\mathrm{A}_{-3}=\{a\}$, $\phi_{-3}=p$, $\mathrm{B}_{0}=\mathtt{AG}$, $\psi_{0}=\bot$, $\mathrm{B}_{1}=\mathtt{AG}$, and $\psi_{1}=p\land q$.
Definition 9 (Modal depth of formulas).
Recursively define the modal depth $\mathsf{md}(\phi)$ of formulas $\phi$ in $\Phi$ as follows:
  • $\mathsf{md}(\top)=(p)=0$;
  • $\mathsf{md}(\neg\psi)=\mathsf{md}(\psi)$;
  • $\mathsf{md}(\psi\land\chi)=\max(\mathsf{md}(\psi),\mathsf{md}(\chi))$;
  • $\mathsf{md}(\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\psi)=\mathsf{md}(\psi)+1$.
@@
Let $\mathsf{X}\in\mathrm{ES}$. For every $\phi\in\Phi$ such that $0<\mathsf{md}(\phi)$, there is $\phi^{\prime}$ such that (1) $\vdash_{\mathsf{MCL}+\mathsf{X}}\phi\leftrightarrow\phi^{\prime}$, (2) $\phi$ and $\phi^{\prime}$ have the same modal depth, and (3) $\phi^{\prime}$ is in the form of $\delta_{0}\land\dots\land\delta_{k}$, where every $\delta_{i}$ is a standard disjunction.
Proof.
Let $\phi\in\Phi$ such that $0<\mathsf{md}(\phi)$.
First, we can transform $\phi$ to $\psi$ in the conjunctive normal form $\psi_{0}\land\dots\land\psi_{k}$ such that $\psi$ has the same modal depth as $\phi$ and $\vdash\phi\leftrightarrow\psi$.
Fix a $\psi_{i}$. Note $\psi_{i}$ is in the form $\alpha_{0}\lor\dots\lor\alpha_{l}$, where every $\alpha_{j}$ is either a propositional literal, or in the form $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\beta$, or in the form $\neg\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\beta$.
Assume there is no $\alpha_{j}$ in the form $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\beta$. Let $\chi_{i}=\psi_{i}\lor\bot\lor\neg\langle\hskip-2.0pt[\mathtt{AG}]\hskip-2.0pt\rangle\bot$. Note $\neg\bot$ and $\neg\langle\hskip-2.0pt[\mathtt{AG}]\hskip-2.0pt\rangle\bot$ are axioms of $\mathsf{MCL}+\mathsf{X}$}. Then $\vdash\psi_{i}\leftrightarrow\chi_{i}$. Assume there is $\alpha_{j}$ in the form $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\beta$. Let $\chi_{i}=\psi_{i}\lor\bot\lor\neg\langle\hskip-2.0pt[\mathtt{AG}]\hskip-2.0pt\rangle\bot\lor\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\top$. Note $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\beta\rightarrow\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\top$ is derivable. Then $\vdash\psi_{i}\leftrightarrow\chi_{i}$. It is easy to see that in the two cases, either $\mathsf{md}(\chi_{i})=\mathsf{md}(\psi_{i})$, or $\mathsf{md}(\chi_{i})=1$.
It is easy to transform $\chi_{i}$ to a standard disjunction $\delta_{i}$ such that $\vdash\chi_{i}\leftrightarrow\delta_{i}$ and $\mathsf{md}(\delta_{i})=\mathsf{md}(\chi_{i})$.
Let $\phi^{\prime}=\delta_{0}\land\dots\land\delta_{k}$. Then $\vdash\phi\leftrightarrow\phi^{\prime}$. It is easy to see that $\phi$ and $\phi^{\prime}$ have the same modal depth.
3.3 Generated submodels
In this subsection, we define generated submodels and state a result about them, which will be used in a result about blueprints in the next subsection.
Definition 10 (Generated submodels).
Let $\mathtt{M}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a model and $s\in\mathtt{ST}$. Define a model $\mathtt{M}^{\prime}=(\mathtt{ST}^{\prime},\mathtt{AC},\mathsf{out}^{\prime}_{\mathtt{AG}},\mathsf{label}^{\prime})$ as follows, called the generated submodel of $\mathtt{M}$ from $s$.
  • $\mathtt{ST}^{\prime}$ is the smallest subset of $\mathtt{ST}$ meeting the following conditions:
    • $s\in\mathtt{ST}^{\prime}$;
    • for every $x\in\mathtt{ST}$, if $x\in\mathsf{out}_{\mathtt{AG}}(s^{\prime},\sigma_{\mathtt{AG}})$ for some $s^{\prime}\in\mathtt{ST}^{\prime}$ and $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, then $x\in\mathtt{ST}^{\prime}$.
  • $\mathsf{out}^{\prime}_{\mathtt{AG}}$ is the restriction of $\mathsf{out}_{\mathtt{AG}}$ to $\mathtt{ST}^{\prime}\times\mathtt{JA}_{\mathtt{AG}}$;
  • $\mathsf{label}^{\prime}$ is the restriction of $\mathsf{label}$ to $\mathtt{ST}^{\prime}$.
The following result indicates that generated submodels preserve truth values.
@@ 3.3 Generated submodels
Let $\mathtt{M}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a model and $s\in\mathtt{ST}$. Let $\mathtt{M}^{\prime}=(\mathtt{ST}^{\prime},\mathtt{AC},\mathsf{out}^{\prime}_{\mathtt{AG}},\mathsf{label}^{\prime})$ be the generated submodel of $\mathtt{M}$ from $s$. Then for every $\phi\in\Phi$, for every $s^{\prime}\in\mathtt{ST}^{\prime}$, $\mathtt{M},s^{\prime}\Vdash\phi$ if and only if $\mathtt{M}^{\prime},s^{\prime}\Vdash\phi$.
This fact is easy to show by induction, and we skip its proof.
3.4 Blueprints and their realization
In this subsection, we define an important notion, blueprints, and prove a result about it. Intuitively, a blueprint can be viewed as a guide for constructing a model. Later in the proof for the downward validity lemma, we will use blueprints to construct the needed models.
@@ 3.4 Blueprints and their realization
A blueprint is a tuple $\mathtt{BP}=(\mathtt{AC},\mathtt{list}_{\mathtt{AG}})$, where:
@@
For any $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, intuitively, $\mathtt{list}_{\mathtt{AG}}(\sigma_{\mathtt{AG}})$ specifies some formulas which $\sigma_{\mathtt{AG}}$ can enable.
We now generalize list functions to all coalitions and define performable joint actions, which have a similar meaning to available joint actions.
Definition 12 (Listing functions for and performable joint actions of coalitions).
Let $\mathtt{BP}=(\mathtt{AC},\mathtt{list}_{\mathtt{AG}})$ be a blueprint.
@@
For every $\mathsf{X}\in\mathrm{ES}$, a blueprint should meet some conditions to be used to construct an $\mathsf{X}$-model, which is what the following definition is about.
@@
Let $\mathsf{X}\in\mathrm{ES}$ and $\mathtt{BP}=(\mathtt{AC},\mathtt{list}_{\mathtt{AG}})$ be a blueprint.
@@
Let $\mathsf{X}\in\mathrm{ES}$ and $\mathtt{BP}=(\mathtt{AC},\mathtt{list}_{\mathtt{AG}})$ be an $\mathsf{X}$-regular blueprint.
@@
Let $\mathsf{X}\in\mathrm{ES}$, $\mathtt{BP}=(\mathtt{AC}_{0},\mathtt{list}_{\mathtt{AG}})$ be an $\mathsf{X}$-regular blueprint, and $\gamma$ be an $\mathsf{X}$-satisfiable elementary conjunction.
We say a pointed $\mathsf{X}$-model $(\mathtt{M},s_{0})$, where $\mathtt{M}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$, realizes $\mathtt{BP}$ and $\gamma$ if the following conditions are met:
@@
Let $\mathtt{BP}=(\mathtt{AC}_{0},\mathtt{list}_{\mathtt{AG}})$ be an $\mathsf{X}$-regular blueprint, and $\gamma$ be an $\mathsf{X}$-satisfiable elementary conjunction.
Construction of a pointed model
Let $\{(\mathtt{M}^{\lambda}_{\phi},s^{\lambda}_{\phi})\mid\lambda\in\mathtt{PJA}_{\mathtt{AG}}\text{ and }\phi\in\mathtt{list}_{\mathtt{AG}}(\lambda)\}$, where $\mathtt{M}^{\lambda}_{\phi}=(\mathtt{ST}^{\lambda}_{\phi},\mathtt{AC}^{\lambda}_{\phi},{\mathsf{out}_{\mathtt{AG}}}^{\lambda}_{\phi},\mathsf{label}^{\lambda}_{\phi})$, be a set of pointed $\mathsf{X}$-models meeting the following conditions:
@@
Let $s_{0}$ be a state not in any of these models. Define a pointed model $(\mathtt{M},s_{0})$, where $\mathtt{M}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$, as follows:
@@
[…]
@@
[…]
The constructed pointed model works
@@
We now show that the last three conditions specified in the definition of realization of blueprints hold.
@@
    Let $s\in\mathsf{out}_{\mathrm{C}}(s_{0},\sigma_{\mathrm{C}})$. Then $s\in\mathsf{out}_{\mathtt{AG}}(s_{0},\sigma_{\mathtt{AG}})$ for some $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$. Then, $s=s^{\sigma_{\mathtt{AG}}}_{\phi}$ for some $\phi\in\mathtt{list}_{\mathtt{AG}}(\sigma_{\mathtt{AG}})$. Note $\mathtt{M}^{\sigma_{\mathtt{AG}}}_{\phi},s^{\sigma_{\mathtt{AG}}}_{\phi}\Vdash\phi$. Then $\mathtt{M},s^{\sigma_{\mathtt{AG}}}_{\phi}\Vdash\phi$. Then $\mathtt{M},s^{\sigma_{\mathtt{AG}}}_{\phi}\Vdash\psi$, that is, $\mathtt{M},s\Vdash\psi$. Therefore, ${\sigma_{\mathrm{C}}}\vartriangleright_{(\mathtt{M},s_{0})}\psi$.